OpenStudy (anonymous):

2a - 15 = -2a + -3 7 - 6a = 6 - 7a 9/2x +5 = 3/2x +17 2(r - 4) = 5(r+ -7) 5a + 6 = -4a + -7 3a - 42 = -3(a+4) 7x -4 = 6x +11 19 - 4a = 3 - 6a

OpenStudy (anonymous):

1)2a+2a=-3+15 4a=12 a=12/4=3 2)7-6a=6-7a -6a+7a=6-7 a=-1 3)9/2x +5 = 3/2x +17 4.5x-1.5x=17-5 3x=12 x=4 4)2(r - 4) = 5(r+ -7) 2r-8=5r-35 -8+35=5r-2r 27=3r 27/3=9=r 5)